# Value Flows

### v1

@@ 2 Related Work
Perhaps the most popular modern generative models are autoencoders (Kingma & Welling, 2013), autoregressive models (Vaswani et al., 2017), denoising diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song et al., 2021), and flow-matching models (Lipman et al., 2023; 2024; Liu et al., 2023; Albergo & Vanden-Eijnden, 2023). In the context of RL, these generative models have succeeded in modeling the trajectories (Chen et al., 2021; Ajay et al., 2023), the transitions (Alonso et al., 2024; Janner et al., 2021; Farebrother et al., 2025), the skills (Ajay et al., 2021; Pertsch et al., 2021; Frans et al., 2024; Zheng et al., 2025), and the policies (Wang et al., 2023b; Hansen-Estruch et al., 2023b; Park et al., 2025c; Dong et al., 2025b). We employ the state-of-the-art flow-matching objective (Lipman et al., 2023) to estimate the future return distribution.
@@ 3 Preliminaries
Flow matching (Lipman et al., 2023; 2024; Liu et al., 2023; Albergo & Vanden-Eijnden, 2023) refers to a family of generative models based on ordinary differential equations (ODEs). The goal of flow matching methods is to transform a simple noise distribution into a target distribution $p_{{\mathcal{X}}}$ over some space ${\mathcal{X}}\subset\mathbb{R}^{d}$. We will use $t$ to denote the flow time step and sample the noise $\epsilon$ from a standard Gaussian distribution ${\mathcal{N}}(0,I)$ throughout our discussions. Flow matching uses a time-dependent vector field $v:[0,1]\times\mathbb{R}^{d}\to\mathbb{R}^{d}$ to construct a time-dependent diffeomorphic flow $\phi:[0,1]\times\mathbb{R}^{d}\to\mathbb{R}^{d}$  (Lipman et al., 2023; 2024) that realizes the transformation. The resulting probability density path $p:[0,1]\times\mathbb{R}^{d}\to\Delta(\mathbb{R}^{d})$ of the vector field $v$ satisfies the continuity equation (Lipman et al., 2023). In Sec. 4, we will estimate the return distribution using flow-based models, utilizing the continuity equation (Sec. 4.2) and flow ODE (Sec. 4.3).
@@ 4.2 Estimating the return distribution using flow-matching
See Appendix B.1 for the detailed statement and a proof. There are two implications of this proposition. First, since the new vector field $v_{k+1}$ generates the new probability density path $p_{k+1}$, applying the distributional Bellman operator to the probability density path $p_{k}$ is equivalent to learning a vector field $v_{k+1}$ that satisfies Eq. 3. This implication allows us to convert the problem of estimating the full return distribution into the problem of learning a vector field, which naturally fits into the flow-matching framework. Second, since the distributional Bellman operator ${\mathcal{T}}^{\pi}$ is a $\gamma$-contraction, repeatly applying ${\mathcal{T}}^{\pi}$ to the probability density path $p_{k}$ guarantees convergence to $p^{\star}(z^{t}\mid t,s,a)$ (Eq. 2). Therefore, we can construct a flow matching loss to learn the vector field $v_{k+1}$ with a guarantee to converge toward $v^{\star}$.
@@
We now derive the preliminary flow matching losses for estimating the return distribution only using transition samples from the dataset DD. [^3] In this case, we do not have full access to the distributions of policy $\pi(a\mid s)$ and transition $p(s^{\prime}\mid s,a)$. Given the vector field $v_{k}$, one simple approach to learn a new vector field $v$ satisfying the update rule in Eq. 3 is minimizing the mean squared error (MSE) between $v$ and $v_{k+1}$. We call this loss the distributional flow matching (DFM) loss:
@@
${\mathcal{L}}_{\text{DCFM}}(v,v_{k})$ $=\mathbb{E}_{\begin{subarray}{c}(s,a,r,s^{\prime})\sim D,t\sim\textsc{Unif}([0,1])\\ a^{\prime}\sim\pi(a^{\prime}\mid s^{\prime}),z^{t}\sim\frac{1}{\gamma}p_{k}\left(\left.\frac{z^{t}-r}{\gamma}\right\rvert t,s^{\prime},a^{\prime}\right)\end{subarray}}\left[\left(v(z^{t}\mid t,s,a)-v_{k}\left(\left.\frac{z^{t}-r}{\gamma}\right\rvert t,s^{\prime},a^{\prime}\right)\right)^{2}\right]$ (5)
Note that we can interpret the transformed returns $(z^{t}-r)/\gamma$ as convolving the probability density path $p_{k}$ and use a change of variable to simplify this DCFM loss (see Lemma 5 in Appendix B). Unlike the DFM loss ${\mathcal{L}}_{\text{DFM}}$, the DCFM loss ${\mathcal{L}}_{\text{DCFM}}$ can be easily estimated via samples from the dataset $D$. Like the connection between the flow matching loss and the conditional flow matching loss in Lipman et al. (2023), the DCFM loss has the same gradient as the DFM loss, indicating that they have the same minimizer.
@@
See Appendix B.1 for the detailed statement and a proof. Importantly, optimizing ${\mathcal{L}}_{\text{DCFM}}$ does not require access to the full transition distribution and prevents taking the intractable integral. Note that ${\mathcal{L}}_{\text{DCFM}}$ is a TD loss because it corresponds to applying the distributional Bellman operator ${\mathcal{T}}^{\pi}$ to the probability density path $p_{k}$. Furthermore, our flow-based model produces estimations of the return expectation (Q function; Sec. 4.4) and the return variance (aleatoric uncertainty; Sec. 4.3) easily, both of which will be incorporated into our algorithm. In Appendix C.1, we discuss the practical flow matching losses for fitting the return distribution.
@@ 4.3 Harnessing uncertainty in the return estimation
One of the benefits of using the flow-based models to estimate the full return distribution is that we can easily compute the return variance at every state-action pair $(s,a)$. These return variances measure the aleatoric uncertainty in the MDP, indicating the noise that stems from the environmental transitions. We incorporate the return uncertainty information into our method and use it to prioritize learning a more accurate return distribution at transitions with higher return variance. We first present an estimation of return expectation using the initial vector field and an approximation of return variance using the Taylor expansion. We then introduce a new ODE that relates the derivative of the learned diffeomorphic flow, allowing us to estimate the return variance in practice. Using this variance estimation, we define the confidence weight for reweighting the distributional flow matching losses.
@@
where $\sigma(\cdot)$ denotes the sigmoid function, $\tau>0$ is a temperature, and $w(s,a,\epsilon)\in[0.5,1]$. For simplicity, we can use one Gaussian noise $\epsilon$ as a Monte Carlo estimator for both the return expectation and the return variance (Eq. 6). We include a visualization of the confidence weight for different temperatures in Appendix E. Our confidence weight is an increasing function with respect to the return variance estimate, indicating that a higher uncertainty results in a higher weight.
@@ 4.4 The complete algorithm
10:  Update the target return vector field $v_{\bar{\theta}}$ using exponential moving averages
@@
Our algorithm, Value Flows, consists of three main components: (1) the vector field estimating the return distribution, (2) the confidence weight incorporating aleatoric uncertainty, and (3) the flow policy selecting actions. Using neural networks to parameterize the return vector field $v_{\theta}$, the BC flow policy $\pi_{\omega}$, and the one-step flow policy $\pi_{\eta}$, we optimize them using stochastic gradient descent. Alg. 1 summarizes the pseudocode of Value Flows, and our open-source implementation is available online. [^4] https://github.com/chongyi-zheng/value-flows
@@ 5 Experiments
The most relevant prior work is distributional RL methods, which model the return distribution using a categorical distribution (C51 (Bellemare et al., 2017)) or a finite number of quantiles (IQN (Dabney et al., 2018) and CODAC (Ma et al., 2021)). We will compare Value Flows against these prior methods by visualizing the return predictions and measuring the success rates in offline and offline-to-online settings. We also compare Value Flows with prior RL methods that learn the scalar Q values with a Gaussian policy (BC (Fujimoto & Gu, 2021), IQL (Kostrikov et al., 2021a), and ReBRAC (Tarasov et al., 2023)) or a flow policy (FBRAC, IFQL, FQL (Park et al., 2025c)) for completeness. Our experiments will use standard benchmarks introduced by prior work on offline RL. We choose a set of 36 state-based tasks from OGBench (Park et al., 2025a) and D4RL (Fu et al., 2021) and choose a set of 25 image-based tasks from OGBench to evaluate our algorithm in the offline setting. Following prior work (Park et al., 2025c), we report means and standard deviations over 8 random seeds for state-based tasks (4 seeds for image-based tasks). We defer the detailed discussions about environments and datasets to Appendix F.1 and baselines to Appendix F.2. We present the hyperparameters in Appendix F.3.
@@ 5.1 Visualizing return distributions of Value Flows
One of the motivations of designing Value Flows is to use flow-based models to estimate the entire return distribution (Sec. 4.1). We hypothesize that the proposed method fits the underlying return distributions with lower distributional discrepancy. To investigate this hypothesis, we study the return predictions from Value Flows and compare them against two alternative distributional RL methods: C51 (Bellemare et al., 2017) and CODAC (Ma et al., 2021). On scene-play-singletask-task2-v0 from OGBench, we visualize the 1D return histograms inferred by different algorithms, including the ground-truth return distribution for reference. For fair comparison, we use $5000$ return samples and $60$ bins for each histogram. For quantitative evaluations, we measure the $1$-Wasserstein distances (Panaretos & Zemel, 2019) between the histogram of each method and the ground-truth histogram.
@@ 5.3 The key components of Value Flows
Figure 5: Reweighing the flow-matching objective boosts success rates. Choosing the correct confidence weight boosts the performance of Value Flows.
Our final set of experiments studies the key components of Value Flows: the regularization loss (Eq. 17), and the confidence weight (Eq. 7). To study the effects of these two components, we conduct ablation experiments on two representative state-based OGBench tasks (cube-double-play-singletask-task2-v0 and puzzle-3x3-play-singletask-task4-v0), varying the regularization coefficient $\lambda$ for the regularization loss and the temperature $\tau$ for the confidence weight.
First, results in Fig. 5 suggest that adding the BCFM regularization loss into the distributional flow matching objective is important for making progress on the puzzle-3x3 task, while choosing the correct regularization coefficient $\lambda$ boosts the performance of Value Flows by $2.6\times$. We also observe that the best $\lambda$ value increases the sample efficiency of our algorithm by $2\times$ compared to the unregularized variant. We conjecture that the BCFM regularization serves as an efficient anchor for learning the full return distribution. Second, Fig. 5 demonstrates that using the confidence weight to reweight the distributional flow matching loss increases the success rate, where an appropriate temperature $\tau$ results in $60\%$ higher success rate on average. In Appendix D.3, we include additional experiments studying the robustness of Value Flows against the number of flow steps (Fig. 8) in the Euler method and the number of candidates in rejection sampling (Fig. 8) for action selections.
@@ 6 Conclusion
Although we focus on learning the full return distribution using flow-matching and reweighting our flow-matching objective using uncertainty estimation. It remains unclear how to disentangle the epistemic uncertainty from the aleatoric uncertainty with the current method. In addition, our discussion is orthogonal to the policy extraction mechanisms. Finding the appropriate policy extraction method within the distributional RL framework might reveal the true benefits of estimating the full return distribution.
@@
We include additional implementation details for hyperparameters, datasets, and evaluation protocols in Appendix F. Our open-source implementations, including baselines, can be found at https://github.com/chongyi-zheng/value-flows. We run experiments for state-based tasks on A6000 GPUs for up to 6 hours, and run experiments for image-based tasks on the same type of GPUs for up to 16 hours.
@@ Acknowledgements
We thank Catherine Ji, Bhavya Agrawalla, and Seohong Park for their helpful discussions. This work used the Della computational cluster provided by Princeton Research Computing, as well as the Ionic and Neuronic computing clusters maintained by the Department of Computer Science at Princeton University. This work was in part supported by an NSF CAREER award, AFOSR YIP, ONR grant N00014-22-1-2293. This material is based upon work supported by the National Science Foundation under Award Numbers 2441665 and 1941722. Any opinions, findings, conclusions, or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the National Science Foundation.

### v2

@@ 2 Related Work
Perhaps the most popular modern generative models are autoencoders (Kingma & Welling, 2013), autoregressive models (Vaswani et al., 2017), denoising diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song et al., 2021), and flow-matching models (Lipman et al., 2023; 2024; Liu et al., 2023; Albergo & Vanden-Eijnden, 2023). In the context of RL, these generative models have succeeded in modeling the trajectories (Chen et al., 2021; Ajay et al., 2023), the transitions (Alonso et al., 2024; Janner et al., 2021; Farebrother et al., 2025b), the skills (Ajay et al., 2021; Pertsch et al., 2021; Frans et al., 2024; Zheng et al., 2025), and the policies (Wang et al., 2023b; Hansen-Estruch et al., 2023b; Dong et al., 2025c; Park et al., 2025c; Dong et al., 2025b). We employ the state-of-the-art flow-matching objective (Lipman et al., 2023) to estimate the future return distribution.
@@ 3 Preliminaries
Flow matching (Lipman et al., 2023; 2024; Liu et al., 2023; Albergo & Vanden-Eijnden, 2023) refers to a family of generative models based on ordinary differential equations (ODEs). The goal of flow matching methods is to transform a simple noise distribution into a target distribution $p_{{\mathcal{X}}}$ over some space ${\mathcal{X}}\subset\mathbb{R}^{d}$. We will use $t$ to denote the flow time step and sample the noise $\epsilon$ from a standard Gaussian distribution ${\mathcal{N}}(0,I)$ throughout our discussions. Flow matching uses a time-dependent vector field $v:[0,1]\times\mathbb{R}^{d}\to\mathbb{R}^{d}$ to construct a time-dependent diffeomorphic flow $\phi:[0,1]\times\mathbb{R}^{d}\to\mathbb{R}^{d}$  (Lipman et al., 2023; 2024) that realizes the transformation. The resulting probability density path $p:[0,1]\times\mathbb{R}^{d}\to\Delta(\mathbb{R}^{d})$ of the vector field $v$ satisfies the continuity equation (Lipman et al., 2023)Appendix A.3 includes the flow ODE (Eq. 11) and the continuity equation (Sec. 12) connecting $\phi$, $v$ and $p$. In Sec. 4, we will estimate the return distribution using flow-based models, utilizing the continuity equation (Sec. 4.2) and flow ODE (Sec. 4.3).
@@ 4.2 Estimating the return distribution using flow-matching
See Appendix B.1 for the detailed statement and a proof. There are two implications of this proposition. First, since the new vector fireld $v_{k+1}$ generates the new probability density path $p_{k+1}$, applying the distributional Bellman operator to the probability density path $p_{k}$ is equivalent to learning a vector field $v_{k+1}$ that satisfies Eq. 3. This implication allows us to convert the problem of estimating the full return distribution into the problem of learning a vector field, which naturally fits into the flow-matching framework. Second, since the distributional Bellman operator ${\mathcal{T}}^{\pi}$ is a $\gamma$-contraction, repeatly applying ${\mathcal{T}}^{\pi}$ to the probability density path $p_{k}$ guarantees convergence to $p^{\star}(z^{t}\mid t,s,a)$ (Eq. 2). Therefore, we can construct a flow matching loss to learn the vector field $v_{k+1}$ with a guarantee to converge toward $v^{\star}$.
@@
We now derive the preliminary flow matching losses for estimating the return distribution only using transition samples from the dataset $D$. Given the vector field $v_{k}$, one simple approach to learn a new vector field $v$ satisfying the update rule in Eq. 3 is to minimize the mean squared error (MSE) between $v$ and $v_{k+1}$. We call this loss the distributional flow matching (DFM) loss:
@@
${\mathcal{L}}_{\text{DCFM}}(v,v_{k})$ $=\mathbb{E}_{\begin{subarray}{c}(s,a,r,s^{\prime})\sim D,t\sim\textsc{Unif}([0,1])\\ a^{\prime}\sim\pi(a^{\prime}\mid s^{\prime}),z^{t}\sim\frac{1}{\gamma}p_{k}\left(\left.\frac{z^{t}-r}{\gamma}\right\rvert t,s^{\prime},a^{\prime}\right)\end{subarray}}\left[\left(v(z^{t}\mid t,s,a)-v_{k}\left(\left.\frac{z^{t}-r}{\gamma}\right\rvert t,s^{\prime},a^{\prime}\right)\right)^{2}\right].$ (5)
We can interpret the transformed returns $(z^{t}-r)/\gamma$ as convolving the probability density path $p_{k}$ and use a change of variable to simplify this DCFM loss (see Lemma 5 in Appendix B). Unlike the DFM loss ${\mathcal{L}}_{\text{DFM}}$, the DCFM loss ${\mathcal{L}}_{\text{DCFM}}$ can be easily estimated via samples from the dataset $D$. Like the connection between the flow matching loss and the conditional flow matching loss in Lipman et al. (2023), the DCFM loss has the same gradient as the DFM loss, indicating that they have the same minimizer.
@@
See Appendix B.1 for the detailed statement and a proof. Importantly, optimizing ${\mathcal{L}}_{\text{DCFM}}$ does not require access to the full transition distribution and prevents taking the intractable integral. ${\mathcal{L}}_{\text{DCFM}}$ is a TD loss because it corresponds to applying the distributional Bellman operator ${\mathcal{T}}^{\pi}$ to the probability density path $p_{k}$. Furthermore, our flow-based model produces estimations of the return expectation (Q function; Sec. 4.4) and the return variance (aleatoric uncertainty; Sec. 4.3) easily, both of which will be incorporated into our algorithm.
We note that simply optimizing ${\mathcal{L}}_{\text{DCFM}}(v,v_{k})$ might produce a degenerated vector field, e.g., $v=0$. In particular, our initial experiments suggest that naively optimizing ${\mathcal{L}}_{\text{DCFM}}(v)$ produced a trivial performance on some tasks. To stabilize learning, we use the return predictions at the next state-action pair $(s^{\prime},a^{\prime})$ and the flow time $t=1$ to construct a bootstrapped target return similar to the TD target in Q-learning (Watkins & Dayan, 1992). The resulting loss function, called bootstrapped conditional flow matching (BCFM) loss ${\mathcal{L}}_{BCFM}$, resembles the standard fitted Q-learning (Riedmiller, 2005). In addition, we use a target vector field $\bar{v}$ to replace the historical vector field $v_{k}$. Prior work has shown that using a target network prevents the model from collapsing (Grill et al., 2020; Tian et al., 2021). See Appendix C.1 for further discussions.
@@ 4.3 Harnessing uncertainty in the return estimation
One benefit of using the flow-based models to estimate the full return distribution is that we can easily compute the return variance at every state-action pair $(s,a)$. These return variances measure the aleatoric uncertainty in the MDP, indicating the noise that stems from the environmental transitions. We incorporate the return uncertainty information into our method and use it to prioritize learning a more accurate return distribution at transitions with higher return variance. We first present an estimation of return expectation using the initial vector field and an approximation of return variance using the Taylor expansion. We then introduce a new ODE that relates the derivative of the learned diffeomorphic flow, allowing us to estimate the return variance in practice. Using this variance estimation, we define the confidence weight for reweighting the distributional flow matching losses.
@@
where $\sigma(\cdot)$ denotes the sigmoid function, $\tau>0$ is a temperature, and $w(s,a,\epsilon)\in[0.5,1]$. For simplicity, we can use one Gaussian noise $\epsilon$ as a Monte Carlo estimator for both the return expectation and the return variance (Eq. 6). We include a visualization of the confidence weight for different temperatures in Appendix D.2. Our confidence weight is an increasing function with respect to the return variance estimate, indicating that a higher uncertainty results in a higher weight. In Appendix C.1, we discuss the complete flow matching losses for fitting the return distribution.
@@ 4.4 The complete algorithm
10:  Update the target return vector field $v_{\bar{\theta}}$ using Polyak averages
@@
Our algorithm, Value Flows, consists of three main components: (1) the vector field estimating the return distribution, (2) the confidence weight prioritizing learning the return distribution at uncertain transition, and (3) the flow policy selecting actions. Using neural networks to parameterize the return vector field $v_{\theta}$, the BC flow policy $\pi_{\omega}$, and the one-step flow policy $\pi_{\eta}$, we optimize them using stochastic gradient descent. Alg. 1 summarizes the pseudocode of Value Flows, and our open-source implementation is available online. [^3] https://github.com/chongyi-zheng/value-flows
@@ 5 Experiments
The most relevant prior work is distributional RL methods, which model the return distribution using a categorical distribution (C51 (Bellemare et al., 2017)) or a finite number of quantiles (IQN (Dabney et al., 2018) and CODAC (Ma et al., 2021)). We will compare Value Flows against these prior methods by visualizing the return predictions and measuring the success rates in offline and offline-to-online settings. We also compare Value Flows with prior RL methods that learn the scalar Q values with a Gaussian policy (BC, IQL (Kostrikov et al., 2021a), and ReBRAC (Tarasov et al., 2023)) or a flow policy (FBRAC, IFQL, FQL (Park et al., 2025c)) for completeness. Our experiments will use standard benchmarks introduced by prior work on offline RL. We choose a set of 36 state-based tasks from OGBench (Park et al., 2025a) and D4RL (Fu et al., 2021) and choose a set of 25 image-based tasks from OGBench to evaluate our algorithm in the offline setting. Following prior work (Park et al., 2025c), we report means and standard deviations over 8 random seeds for state-based tasks (4 seeds for image-based tasks). We defer the detailed discussions about environments and datasets to Appendix E.1 and baselines to Appendix E.2. We present the hyperparameters in Appendix E.3.
@@ 5.1 Visualizing return distributions of Value Flows
One of the motivations of designing Value Flows is to use flow-based models to estimate the entire return distribution (Sec. 4.1). We hypothesize that the proposed method fits the underlying return distributions with lower distributional discrepancy. To investigate this hypothesis, we study the return predictions from Value Flows and compare them against two alternative distributional RL methods: C51 (Bellemare et al., 2017) and CODAC (Ma et al., 2021). On scene-play-singletask-task2-v0 from OGBench, we use a learned FQL agent to collect $5000$ optimal and suboptimal trajectories as our evaluation dataset to visualize the return distribution for different methods. These trajectories are out-of-distribution because (1) they are collected by a learned agent instead of the behavioral policy, and (2) they yield some combinatorial generalization behaviors (see Appendix E.1 for explanations). We visualize the 1D return histograms inferred by different algorithms, including the ground-truth return distribution for reference. For fair comparison, we use $5000$ return samples and $60$ bins for each histogram. For quantitative evaluations, we measure the $1$-Wasserstein distances (Panaretos & Zemel, 2019) between the histogram of each method and the ground-truth histogram.
@@ 5.3 The key components of Value Flows
Our final set of experiments studies the key components of Value Flows: the regularization loss (Eq. 17), and the confidence weight (Eq. 7). To study the effects of these two components, we conduct ablation experiments on two representative state-based OGBench tasks, varying the regularization coefficient $\lambda$ for the regularization loss and the temperature $\tau$ for the confidence weight.
@@
Figure 5: Reweighting the flow-matching objective boosts success rates. Choosing the correct confidence weight boosts the performance of Value Flows.
First, results in Fig. 5 suggest that adding the BCFM regularization loss into the distributional flow matching objective is important for making progress on the puzzle-3x3 task, while choosing the correct regularization coefficient $\lambda$ boosts the performance of Value Flows by $2.6\times$. We also observe that the best $\lambda$ value increases the sample efficiency of our algorithm by $2\times$ compared to the unregularized variant. We conjecture that the BCFM regularization serves as an efficient anchor for learning the full return distribution. Second, Fig. 5 demonstrates that using the confidence weight to reweight the distributional flow matching loss increases the success rate, where an appropriate temperature $\tau$ results in a $60\%$ higher success rate on average. In Appendix D.1, we include additional experiments studying the robustness of Value Flows against the number of flow steps (Fig. 8) in the Euler method and the number of candidates in rejection sampling (Fig. 8) for action selections.
Additional experiments.
In Appendix D.3, we investigate the computational time and GPU memory consumption of Value FlowsAppendix D.4 demonstrates the capability of Value Flows in an MDP requiring exploration and risk-sensitive control. In Appendix D.5, we include an ablation study showing that using Gaussian policies improves the performance of Value Flows on D4RL adroit tasks. The ablations in Appendix D.6 study the role of the ${\mathcal{L}}_{\text{BCFM}}$ loss, which mainly serves as a regularization.
@@ 6 Conclusion
Although we focus on learning the full return distribution using flow-matching and reweighting our flow-matching objective using uncertainty estimation, it remains unclear how to disentangle the epistemic uncertainty from the aleatoric uncertainty with the current method. In addition, our discussion is orthogonal to the policy extraction mechanisms. Finding the appropriate policy extraction method within the distributional RL framework might reveal the true benefits of estimating the full return distribution.
@@
We include additional implementation details for hyperparameters, datasets, and evaluation protocols in Appendix E. Our open-source implementations, including baselines, can be found at https://github.com/chongyi-zheng/value-flows. We run experiments for state-based tasks on A6000 GPUs for up to 6 hours, and run experiments for image-based tasks on the same type of GPUs for up to 16 hours.
@@ Acknowledgements
We thank Catherine Ji, Bhavya Agrawalla, and Seohong Park for their helpful discussions. We thank Seohong Park for sharing the data for some baselines in our experiments. This work used the Della computational cluster provided by Princeton Research Computing, as well as the Ionic and Neuronic computing clusters maintained by the Department of Computer Science at Princeton University. This work was in part supported by an NSF CAREER award, AFOSR YIP, ONR grant N00014-22-1-2293. This material is based upon work supported by the National Science Foundation under Award Numbers 2441665, 2237693, and 1941722. Any opinions, findings, conclusions, or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the National Science Foundation.